OpenStudy (anonymous):

Simplify: iÖ-98 - Ö98

OpenStudy (anonymous):

What was that?? Ha ha ha..

OpenStudy (anonymous):

Ö what is this?

OpenStudy (anonymous):

i√-98 - √98

OpenStudy (anonymous):

Square root symbol

OpenStudy (anonymous):

√-98=i√98 use this

OpenStudy (anonymous):

Use this also: \[\sqrt{98} = 7\sqrt{2}\]

OpenStudy (anonymous):

so it becomes: i(i√98) -√98=-2√98=-14√2

OpenStudy (anonymous):

remmember that i*i=-1